Required information

The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,750 $ 16,650 $ 29,400
Estimated variable manufacturing overhead per machine-hour $ 2.50 $ 3.30
Job P Job Q
Direct materials $ 24,000 $ 13,500
Direct labor cost $ 29,800 $ 11,900
Actual machine-hours used:
Molding 2,800 1,900
Fabrication 1,700 2,000
Total 4,500 3,900

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-9, assume that Sweeten Company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments and Job P included 20 units and Job Q included 30 units. For questions 10-15, assume that the company uses a plantwide predetermined overhead rate with machine-hours as the allocation base.

4. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.)

5. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

6. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

7. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

8. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

9. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

10. What was the company’s plantwide predetermined overhead rate? (Round your answer to 2 decimal places.)

11. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

Molding overhead rate = (12750/2500+2.5) = 7.6 per hour

Fabrication overhead rate = (16650/1500+3.3) = 14.40 per hour

4) Total manufacturing cost

Job P
Direct material 24000
Direct labor 29800
Molding overhead (2800*7.6) 21280
Fabrication overhead (1700*14.40) 24480
Total manufacturing cost 99560

5) Unit product cost = 99560/20 = 4978

6) Total manufacturing cost

Job Q
Direct material 13500
Direct labor 11900
Molding overhead (1900*7.6) 14440
Fabrication overhead (2000*14.40) 28800
Total manufacturing cost 68640

7) Unit product cost = 68640/30 = 2288

Note: Pls post whole que in the group of 4-4 sub parts as per HOMEWORKLIB RULES